OpenStudy (anonymous):

Can u solve: A=P(1+r/n) with nt exponent when P=5000, r=.05, n=12, t=8 I can get as far as appling the exponent to find the answer, then I am stuck.

OpenStudy (anonymous):

A = 5000 (1 + 0.05/12) ^(12*8) A= 5000(241/240)^96 A = 5000(1.4906) A =7,452.93